OpenStudy (anonymous):

does 2pi * 4pi = 6pi or 6pi^2?

OpenStudy (anonymous):

8pi^2

OpenStudy (kinggeorge):

You need to also multiply the numbers instead of adding them. So sandbox is correct. It should be \(8\pi^2\)

OpenStudy (callisto):

2pi * 4pi = 2*4 *pi *pi = 8pi^2